OpenStudy (loser66):

Please, give me an example to illustrate this expression T is linear transformation on C Tz = az + comples conjugate (z)

OpenStudy (hba):

why is everyone doing linear algebra today :/

OpenStudy (loser66):

@ikram002p if \(z = \alpha x + \beta i y \), then Tz = \(a(\alpha x+ \beta iy -\beta iy = a\alpha x\) @hba why not?

OpenStudy (loser66):

I have to prove/disprove it is invertible but not sure how to translate the question.

OpenStudy (loser66):

no idea?

OpenStudy (ikram002p):

sry was afk , ill check this :D

OpenStudy (loser66):

I appreciate your quickly response whenever I tag. :)

OpenStudy (ikram002p):

when z= ax+Biy conjugate (z)=ax-Biy so ur exampl Tz=a(ax+Biy)+(ax-Biy) Tz=(a^2+a)x + (aB-B) y

OpenStudy (ikram002p):

so , ok x , y still in linear mood a,b constant right ? then it work

OpenStudy (ikram002p):

but need to make sure that Tz is not = aax in the example u gave

OpenStudy (loser66):

I don't think a in Tz =a z is a in z = ax + bi y that's why I put z = \(\alpha x +\beta iy\) to make them separate

OpenStudy (ikram002p):

i see so u have a and alpha ?

OpenStudy (ikram002p):

well , it wont change it well still linear cuz x,y both linear exchange

OpenStudy (loser66):

Thanks for your explanation. I think I can work now. hihihi

OpenStudy (ikram002p):

good ! gtg now , good luck with the rest of it :D

OpenStudy (loser66):

:)